RELATION.

Given the following set:

1. X = {1, 2, 3, 4, 5} defined by the rule (x, y) ∈ R if x + y ≤ 6 


a. List the elements of R

b. Find the domain of R

c. Find the range of R

d. Draw the digraph

e. Properties of the Relation 


1
Expert's answer
2021-03-16T08:19:25-0400

Given the "X = \{1, 2, 3, 4, 5\}" define the relation "R" by the rule "(x, y) \in R" if "x + y \le 6"  


a. Let us find the list the elements of "R:"


"R=\{(1,1),(1,2),(2,1),(1,3),(3,1),(2,2),(1,4),(4,1),(2,3),(3,2),(1,5), \newline (5,1),(2,4),(4,2),(3,3)\}"


b. Let us find the domain of "R:"


"dom(R)=\{x\in X\ |\ (x,y)\in R\}=\{1,2,3,4,5\}"



c. Let us find the range of "R:"


"range(R)=\{y\in X\ |\ (x,y)\in R\}=\{1,2,3,4,5\}"


d. Let us draw the digraph of "R:"




e. Let us study some properties of the relation "R:"


Since "(4,4)\notin R", the relation "R" is not reflexive.


If "(x, y) \in R", then "x + y \le 6", and hence "y + x \le 6". It follows that "(y, x) \in R", and the relation is symmetric.


Taking into account that "(5,1)\in R" and "(1,5)\in R", but "(5,5)\notin R", we conclude that "R" is not a transitive relation.


It follows that "R" is not an equivalence relation.